# µnit Scaling: Simple and Scalable FP8 LLM Training

### v2

@@ 3.3 FP8 Training Efficiency
To achieve state-of-the-art FP8 distributed training efficiency with µnit Scaling, we make use of operator fusion and static scaling. As shown in Fig. 8, FP8 training with µS is 25-33% faster than in BF16, and 1-6% faster than FP8 training with TransformerEngine (TE) NVIDIA (2023). All models were benchmarked on 64 NVIDIA H100 GPUs, and characteristics such as batch size and distributed training configuration were held constant. While TransformerEngine has fused modules such as LayerNorm-Linear or LayerNorm-MLP, we did not use those modules in order to make an equal comparison between µS and TE.
@@
By relying on dynamic scaling, FP8 training with libraries like TE imposes addtional overhead that is eliminated in µS. Calculating the absolute max of both the weight and activation tensors (or storing and reading past absolute max values in a delayed scaling approach) are operations that can be completely discarded in µS. Weights, activations, and gradients can be directly cast to FP8 formats, with a constant $\alpha=\frac{1}{\sqrt{fan\_in}}$ scaling factor used in the hidden linear layers’ GEMM calls, where a GEMM is defined as:
@@
NVIDIA’s H100 GPUs support FP8 GEMMs through the $cublasLtMatmul()$ operation NVIDIA Corporation (2024).
To maximize training speed and mirror TransformerEngine NVIDIA (2023), we fuse clipping to the FP8 range, casting to FP8, and transposing into a single Triton Tillet et al. (2019) kernel. A transpose is necessary because H100s only support one layout (“TN”) with FP8, but the forward and backward passes use different layouts (thanks to using $\mathbf{W}$ vs $\mathbf{W}^{T}$).
@@ Appendix A Appendix
A.1 Covariance of softmax numerator and denominator
@@
A.2 Modifying Residual Connections with $\tau$
To make skip connections variance-preserving, we use the fixed residual modification scheme, as shown in Eq. 10, with coefficients based on the hyperparameter $\tau$ Blake et al. (2023). To understand the relationship of the optimal residual coefficient $\tau^{*}$ with network depth, we swept over various values of $\tau$ for models of different widths (256, 512, 1024, 2048) and depths (20, 40, 60, 80, 100). In order to assess potential confounding effects between $\tau^{*}$ and $\eta^{*}$ and $\lambda^{*}$, we tuned those two hyperparameters as well. We trained each model for 10.5B tokens with a global batch size of 256 and sequence length of 4096. We define the optimal subset of models as those which had a final cross-entropy loss within 0.25% of the optimum (with loss averaged over the last 10 steps, i.e. 10.5M tokens). As shown in Fig. 9, $\tau^{*}$ (for the optimal subset of models) decreases as network depth increases. Since the contribution of each residual branch exponentially decays with depth, a lower $\tau$ ensures a lower rate of decay, likely useful as networks get deeper. This relationship between $\tau^{*}$ and depth is consistent even as model width increases. In our experiments, $\tau$ can be coarsely swept. We use the results shown in Fig. 9, to directly choose $\tau^{*}$ for all µS model training.
@@ A.2 Modifying Residual Connections with $\tau$
A.3 Lion Optimizer and Hyperparameter Transfer
Here, we show why Lion Chen et al. (2023) is an ”Adam-like” optimizer, so the µP rules for hyperparameter transfer with Adam Kingma & Ba (2017) are applicable to Lion as well. Because Adam and Lion are both adaptive optimizers that normalize gradients coordinatewise before updating parameters, the nonlinear tensor product matrix results obtained in Yang et al. (2021, Appendix J.1.3) apply to both optimizers. One can see that Lion differs from Adam only in that it has a different second moment estimate. Under both optimizers, with gradient $g_{t}$, a parameter $\theta$ is updated as:
@@ A.3 Lion Optimizer and Hyperparameter Transfer
A.4 Activation Outliers
@@
A.5 Activation Function Choice
The choice of activation function can have a signficant impact on activation underflow when training in FP8. For example, recent work by Anonymous (2024) identifies outlier amplification from SwiGLU as a challenge for FP8 LLM training. Nearly all state-of-the-art LLMs today use either SiLU or GELU as their activation function, but when training in FP8, this may lead to underflow in activations during training. This is because these functions asymptotically approach zero as inputs $x\rightarrow-\infty$. We define the FP8 underflow fraction, or the fraction of elements flushed to 0 from a BF16 to FP8 cast, as a metric to evaluate various activation functions. As shown in Fig. 10, this can cause many activations to underflow.
@@ A.5 Activation Function Choice
To better understand how activation function choice influences FP8 underflow when training with µnit scaling, we train small 4 layer models with GELU, SiLU, and ReLU. Our findings, detailed in Fig. 11 that during unit scaled model training, the choice of activation function drastically impacts the FP8 underflow rate for activation outputs. GELU greatly degrades the representation of FFN down projection inputs, reaching up to 30% underflow during training. SiLU causes similar degradation, but at a lower rate, reaching up to 7% during training. In contrast, ReLU does not suffer from this problem, with a maximum of 0.04% FP8 underflow during training. As a result, FP8 unit scaled models trained with ReLU have smaller low-precision convergence error (defined as the percent difference between the final cross entropy loss an FP8 model and its BF16 counterpart). Based on these observations and results, ReLU minimizes FP8 underflow and low-precision convergence error. ReLU also has the added benefit of sparsifying activations, a property which enables significant inference-time optimizations Mirzadeh et al. (2024). However, using GELU results in models with lower final training loss. For this reason, we use GELU when training all µS models. Additional investigations into activation functions more suitable for FP8 training can help mitigate underflow while also providing improved convergence.

### v3

@@ 3.3 FP8 Training Efficiency
To achieve state-of-the-art FP8 distributed training efficiency with µnit Scaling, we make use of operator fusion and static scaling. As shown in Fig. 8, FP8 training with µS is 25-33% faster than in BF16, and 1-6% faster than FP8 training with TransformerEngine (TE) (NVIDIA, 2023). All models were benchmarked on 64 NVIDIA H100 GPUs, and characteristics such as batch size and distributed training configuration were held constant. While TransformerEngine has fused modules such as LayerNorm-Linear or LayerNorm-MLP, we did not use those modules in order to make an equal comparison between µS and TE.
@@
By relying on dynamic scaling, FP8 training with libraries like TE imposes additional overhead that is eliminated in µS. Calculating the absolute max of both the weight and activation tensors (or storing and reading past absolute max values in a delayed scaling approach) are operations that can be completely discarded in µS. Weights, activations, and gradients can be directly cast to FP8 formats, with a constant $\alpha=\frac{1}{\sqrt{\text{fan\_in}}}$ scaling factor used in the hidden linear layers’ GEMM calls, where a GEMM is defined as:
@@
NVIDIA’s H100 GPUs support FP8 GEMMs through the $cublasLtMatmul()$ operation (NVIDIA Corporation, 2024).
To maximize training speed and mirror TransformerEngine NVIDIA (2023), we fuse clipping to the FP8 range, casting to FP8, and transposing into a single Triton (Tillet et al., 2019) kernel. A transpose is necessary because H100s only support one layout (“TN”) with FP8, but the forward and backward passes use different layouts (thanks to using $\mathbf{W}$ vs $\mathbf{W}^{T}$).
@@ Appendix A Appendix
A.1 Why these modifications?
Table 1 contains a number of modifications to standard bf16 training setups. Where did these come from? Are they simply a result of trying ideas until something worked? Or are they the result of more principled analysis and ablations?
While we do explain the basis for each modification over the course of the main text, this section summarizes how we arrived at each of them. We can group the origins of these changes into three categories: simple math, adhering to prior art, and ablation experiments.
A.1.1 Simple math
Recall that, in order to ensure stable training and consistent hyperparameter meanings, we wish to ensure that all weight and activation tensors have unit variance. Enforcing unit variance is difficult because the weights are constantly being modified throughout training. To enforce exact unit variance everywhere would require significant overhead in the form of added normalization operations. We therefore relax the constraint to the following:
  1. 1.
    Each residual branch must have exactly unit variance
  2. 2.
    Weight tensors must have unit variance at initialization
  3. 3.
    Linear layer outputs have unit variance at initialization, assuming the inputs are iid with unit variance.
  4. 4.
    Weight updates should attempt to preserve the weight and activation variances to the extent that this is possible without significant overhead.
The last three requirements mirror Blake et al. (2023) while the first is stronger.
Our core modifications follow immediately from these requirements and a bit of math.
[…]
A.1.2 Adhering to best practices.
Some aspects of our training recipe are crucial but already common (though not universal) practices. These include:
Weight decay ($\lambda$) scaling. Recall that decoupled weight decay amounts to multiplying weights by a constant $11-\lambda,0<=\lambda<$ 1 during each update. This operation already has the same semantics across model widths.
FP8 hidden layers. Using e4m3 weights and activations along with e5m2 gradients is a common practice (NVIDIA, 2023; Micikevicius et al., 2022) Clipping instead of overflowing prevents NaN/Inf values. Keeping the first and last layers in higher precision is also common.
A.1.3 Ablation experiments.
Two modifications in our recipe can be implemented in multiple ways, so we chose the details based on smaller-scale experimental results.
Fixed residual modification. In order to satisfy our design goal of having a fixed-variance residual stream, we need to combine the previous residual stream tensor and the latest residual branch output in some manner that preserves variance. As discussed in the paper, this can be done by replacing summation with weighted summation. However, we are left with a degree of freedom in setting the weighting coefficient. To keep the search space small, we consider only the two schemes from (Blake et al., 2023) and decide between them based on the experiments in Section A.3.
Res-Post-LayerNorm. As we show in Section 2.1, the variance of token representations tends to collapse later in the sequence. If a closed-form correction could exactly undo this effect, we could apply such a correction and avoid modifying the architecture. However, as shown in Figures 2 and 3, the pattern of variance collapse is input-dependent and deviates greatly from what iid assumptions would lead one to expect. In order to satisfy our requirement that residual streams have unit variance, we therefore must resort to a blunt instrument: imposing normalization at runtime. We could normalize the residual stream itself, add a normalization op at the end of each residual branch, or move the normalization in a Pre-LN transformer from the start of the branch to the end. We decided to go with the last option because it adds no extra operations, normalizes both the residual stream token embeddings and their updates, is consistent with previous work (Liu et al., 2022; OLMo et al., 2024), and worked well in our ablation experiments (Fig 4(b)).
A.1.4 Comparison to existing schemes
As a supplement to to Table 1 which enumerates the components of µS compared to standard practice (SP), Table 6 compares these components with µP, Unit Scaling, and u-µP.
Table 6: Comparing µS with other schemes µS components have commonalities and differences with existing training schemes. It is the only one which combines scalable, complete FP8 LLM training with hyperparameter transfer; see Figure 1 for a comparison of features of low-precision training methods.
µS Component µP Unit Scaling u-µP
Linear layer scaling factors Not used Used, but can be different in forward and backward pass. Used
Res-Post-LayerNorm Not used Not used Not used
“Fixed” residual modification Not used Proposed Not used
Unit variance initialization Not used Used Used
FP8 hidden layers Not used Used, but not at scale Used, but restricted only to some layers
Learning rate ($\eta$) scaling Used Not used Used
Weight decay ($\lambda$) scaling Not used Not used Used
A.2 Covariance of softmax numerator and denominator
@@
A.3 Modifying Residual Connections with $\tau$
To make skip connections variance-preserving, we use the fixed residual modification scheme, as shown in Eq. 11, with coefficients based on the hyperparameter $\tau$ (Blake et al., 2023). To understand the relationship of the optimal residual coefficient $\tau^{*}$ with network depth, we swept over various values of $\tau$ for models of different widths (256, 512, 1024, 2048) and depths (20, 40, 60, 80, 100). In order to assess potential confounding effects between $\tau^{*}$ and $\eta^{*}$ and $\lambda^{*}$, we tuned those two hyperparameters as well. We trained each model for 10.5B tokens with a global batch size of 256 and sequence length of 4096. We define the optimal subset of models as those which had a final cross-entropy loss within 0.25% of the optimum (with loss averaged over the last 10 steps, i.e. 10.5M tokens). As shown in Fig. 9, $\tau^{*}$ (for the optimal subset of models) decreases as network depth increases. Since the contribution of each residual branch exponentially decays with depth, a lower $\tau$ ensures a lower rate of decay, likely useful as networks get deeper. This relationship between $\tau^{*}$ and depth is consistent even as model width increases. In our experiments, $\tau$ can be coarsely swept. We use the results shown in Fig. 9, to directly choose $\tau^{*}$ for all µS model training.
@@ A.3 Modifying Residual Connections with $\tau$
A.4 Lion Optimizer and Hyperparameter Transfer
Here, we show why Lion Chen et al. (2023) is an ”Adam-like” optimizer, so the µP rules for hyperparameter transfer with Adam (Kingma & Ba, 2017) are applicable to Lion as well. Because Adam and Lion are both adaptive optimizers that normalize gradients coordinatewise before updating parameters, the nonlinear tensor product matrix results obtained in Yang et al. (2021, Appendix J.1.3) apply to both optimizers. One can see that Lion differs from Adam only in that it has a different second moment estimate. Under both optimizers, with gradient $g_{t}$, a parameter $\theta$ is updated as:
@@ A.4 Lion Optimizer and Hyperparameter Transfer
A.5 µnit Scaling vs Unit Scaling for larger model training
We test the unit scaling (US) and µnit scaling (µS) methods at the 7B model scale with FP8 training. Figure 10 shows that unit scaling models diverge very early in training, while µnit scaling runs converge smoothly. Based on this experiment, we did not conduct final model runs at different model scales with unit scaling (1B–13B).
Figure 10: Unit Scaling (US) vs µnit Scaling (µS) for 7B models. Convergence test loss curves at 7B model scale show that µS converges smoothly while US training diverges early in training.
A.6 Activation Outliers
@@
A.7 Activation Function Choice
The choice of activation function can have a significant impact on activation underflow when training in FP8. For example, recent work by (Anonymous, 2024) identifies outlier amplification from SwiGLU as a challenge for FP8 LLM training. Nearly all state-of-the-art LLMs today use either SiLU or GELU as their activation function, but when training in FP8, this may lead to underflow in activations during training. This is because these functions asymptotically approach zero as inputs $x\rightarrow-\infty$. We define the FP8 underflow fraction, or the fraction of elements flushed to 0 from a BF16 to FP8 cast, as a metric to evaluate various activation functions. As shown in Fig. 11, this can cause many activations to underflow.
@@ A.7 Activation Function Choice
To better understand how activation function choice influences FP8 underflow when training with µnit scaling, we train small 4 layer models with GELU, SiLU, and ReLU. Our findings, detailed in Fig. 12 that during unit scaled model training, the choice of activation function drastically impacts the FP8 underflow rate for activation outputs. GELU greatly degrades the representation of FFN down projection inputs, reaching up to 30% underflow during training. SiLU causes similar degradation, but at a lower rate, reaching up to 7% during training. In contrast, ReLU does not suffer from this problem, with a maximum of 0.04% FP8 underflow during training. As a result, FP8 unit scaled models trained with ReLU have smaller low-precision convergence error (defined as the percent difference between the final cross entropy loss an FP8 model and its BF16 counterpart). Based on these observations and results, ReLU minimizes FP8 underflow and low-precision convergence error. ReLU also has the added benefit of sparsifying activations, a property which enables significant inference-time optimizations (Mirzadeh et al., 2024). However, using GELU results in models with lower final training loss. For this reason, we use GELU when training all µS models. Additional investigations into activation functions more suitable for FP8 training can help mitigate underflow while also providing improved convergence.